OpenStudy (anonymous):

What is the greatest possible percent error in calculating the volume of a box measured to the nearest inch with sides of 5 in., 10 in., and 10 in.? Round to the nearest percent. A. 18% B. 21% C. 23% D. 25%

OpenStudy (unklerhaukus):

\[ { \Delta V \over V } = \sqrt {1/ \Delta x_1^2 + 1/\Delta x_2^2 +1/\Delta x_3^2} \]\[= \sqrt {1/5^2 + 1/10^2 +1/10^2} \]\[ = \sqrt {1/25 + 1/100 +1/100} \]\[=\sqrt {6/100}= 0.2449 \]= 25%

OpenStudy (anonymous):

wrong

OpenStudy (unklerhaukus):

explain @faithstarle